Problem 7-05 Nonconstant Growth Valuation A company currently pays a dividend of $3.5 per share (D0 = $3.5). It is estimated that the company's dividend will grow at a rate of 21% per year for the next 2 years, and then at a constant rate of 8% thereafter. The company's stock has a beta of 2, the risk-free rate is 4.5%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=4.5+(2*3)=10.5%

D1=(3.5*1.21)=4.235

D2=(4.235*1.21)=5.12435

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(5.12435*1.08)/(0.105-0.08)

=221.37192

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.235/1.105+5.12435/1.105^2+221.37192/1.105^2

=$189.33(Approx).